NvidiaIntely (nvidiaintely):

Challenge: The distance around the middle of a sphere (the circumference) equals 37.699 in. What is the surface area of the sphere?

NvidiaIntely (nvidiaintely):

@NerdyChick16

OpenStudy (nerdychick16):

"the distance around the middle" means the circumference of the sphere, which is a circle. use circumference of circle formula C = 2* pi* r = 37.699" Now use surface are of sphere formula = 4* pi* r^2 this is the same as (2* pi* r) x (2* pi* r) / pi ..... yes ?!? :) And we know that (2* pi* r) is C !! So surface area of a sphere, if you KNOW the circumference = C* C / pi So the area is = 37.699^2 / pi = 452.39 sq in https://answers.yahoo.com/question/index?qid=20100505082645AAs5LXI

OpenStudy (nerdychick16):

did that help?

NvidiaIntely (nvidiaintely):

Thanks!